Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
vardaro, zendri , zambon - 21:12 Friday 20 February 2026 (68726) Print this report
Recovery of the alignment between SQZ and ITF

We started the shift opening the SDB1 shutter at 8:42UTC with +41000 steps

In the mean time we also checked the status of the squeezer with the HD detector: +8.7 dB of ASQZ and 6dB of SQZ at about 9:10 UTC

After that we opened SQB1 shutter with -500000 steps at 9:48 UTC we struggled to see any beam from ITF

We then understood that the SQB1 AA mirrors were not in the position of last time when we injected SQZ. I.e. 2nd september 2025. We then put them in their last position and this allowed us to see a beam into the SQB1 cameras

2026-02-20-10h24m00-UTC    info vardaro      'AA Mirror1:X ramp [Frequency=0.5,ampl=0,bias=5.28]' sent to SQZ_PLL_AA
2026-02-20-10h24m33-UTC    info vardaro      'AA Mirror1:Y ramp [Frequency=0.5,ampl=0,bias=-1.795]' sent to SQZ_PLL_AA
2026-02-20-10h25m20-UTC    info vardaro      'AA Mirror2:X ramp [Frequency=0.5,ampl=0,bias=0.606]' sent to SQZ_PLL_AA
2026-02-20-10h25m40-UTC    info vardaro      'AA Mirror2:Y ramp [Frequency=0.5,ampl=0,bias=3.22]' sent to SQZ_PLL_AA

 

Once the ITF was locked in LN3 aligned we started the realigment of the squeezer. At 10:40UTC we started the alignment with a magnitude of about 0.1mV or less. The maximum found in 2024 was about 2 mV.

We started to check the CC offset but this was ok

We inserted some ramp on SQB1 mirrors but this was not useful

We then decided to engage the SQZ AA system and we left it drifting. With that we managed to arrive at about 0.8 mV

We started to move SQB1 X and TX direction. We improved a bit going with SQB1 in TX from 300 to 120 urad

We arrived with a B1 4MHz magnitude of about 1.7mV and we injected SQZ into the ITF during lunch break at 12:42:30 UTC. Probably a bit less than one dB was visible in the ITF (fig 1)

Then we started to align with AA and pico motors on SQB1 in order to improve the magnitude. And after some iteration between SQB1 bench, SQZ AA actuators and SQB1 picomotors we mamaged to have a magnitude of about 2.6mV

But with the SQB1 AA mirrors near saturation SQB1 M1 X at +7V and SQB1 M1 Y at -8V, whereas M2 X was at -5.5V and M2 Y of about 0.5 V. Then the interferometer unlocked,

 

ITF relocked at 16 UTC, we waited up to 16:13 to open the shutter then we restarted with the alignment in order to discharge the AA actuators, We tried in closed loop, then the interferometer immediately unlocked.

 

ITF relocked around 17UTC but it ulocked quite immediately after the restart of the alignment

When it relocked the magnitude was very good and the SQB1 actuators far from the saturation. So we put in SQZ from 17:44:27 UTC. Fig2

Shot noise reference at 18:02:00 UTC

Phase scan run at 18:09 UTC Gain 50000 Shot noise reference 2 mins. (fig3)

The phase scan was too short to see the minimum of the magnitude. So we run onother one at 19:17 UTC between 80 and 290 deg (fig4)

After the scan we took 10 min of SQZ at 19:48:24 UTC with an angle of 4.2 rad 

We stopped at 20:02 UTC

Final position of the various DOFs:

  • SQB1 M1 X 5.8V
  • SQB1 M1 Y -4.99V
  • SQB1 M2 X -4.94 V
  • SQB1 M2 Y 1.78V
  • SQB1 TX 0urad
  • SQB1 TY -140urad
  • SQB1 TZ 125 urad
  • SQB1 X -3250 um
  • SQB1 Y -550 um
  • SQB1 Z 710 um 

At the end we closed only SQB1 shutter, we switched off all the dither lines and we rotated HWP2 of SQB1 to send SQZ toward the HD detector

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×